# A Plug-and-Play Method for Guided Multi-contrast MRI Reconstruction based on Content/Style Modeling

### v3

@@ 4.2 Implementation details
We used the same general residual architecture for MUNIT encoders and decoders as the original paper, except with an additional layer at the end of the content encoder to produce content maps of specifies number of channels. We used 2 or 4 channels, depending on the content downsampling factor. The discriminators were implemented as multi-scale PatchGAN networks enabling them to locally assess the input images for realism at different scales. We used 1 scale in the BrainWeb simulations and 3 scales in the NYU and LUMC benchmark. We additionally conditioned the discriminators on foreground masks of the images to penalize background signal. To stabilize GAN training, we used spectral normalization (Miyato et al., 2018). For hyperparameters, we found that pre-training loss weight values and paired fine-tuning values worked generally well for the in vivo datasets. In the benchmarks, the step size of the CR module of PnP-CoSMo was tuned per acceleration for each dataset and the number of iterations was set to 200.
In the simulations, all content/style models were pre-trained for 200k and were not fine-tuned for the sake of simplicity, unless stated otherwise in Section 5.1. In both NYU and LUMC benchmarks, the content/style models were pre-trained for 400k iterations and fine-tuned for 50k iterations. All software was implemented in PyTorch, and image registration was performed using Elastix (Klein et al., 2009). All training runs were performed on a compute node with an NVIDIA Quadro RTX 6000 GPU.
@@ 5 Results
5.1 Experiments on simulated MR datasets
@@ 5.1.2 Content capacity analysis
Table 4 compares PnP-CoSMo reconstruction quality across these configurations. PFT and CR were disabled here for simplicity. We observe two trends. First, the reconstruction quality generally decreases from left to right, eventually dropping below the conventional L1-wavelet CS reconstruction, suggesting a decrease in the amount of shared local information contained in the reference contrast. Second, the optimal content capacity of the model decreases with the reference resolution in accordance with the actual amount of this shared information. In other words, this optimal content capacity corresponds to the amount of contrast-independent structure the model discovers in the dataset, e.g. RefRes-1 dataset had 4 times as much shared structure (due to twice as large optimal content resolution) as RefRes-1/2 and RefRes-1/4 datasets and hence, the RefRes-1 dataset was significantly more effective in guided reconstruction. On the other hand, RefRes-1/2 and RefRes-1/4 datasets had similar levels of shared structure (i.e. =1/22=1/4 times the full-resolition case) despite the lower reference resolution in RefRes-1/4, suggesting a more complex relationship between the reference resolution and the shared content. In the RefRes-1/8 case, PnP-CoSMo dropped to a similar level of quality as conventional CS, suggesting a lowerbound on the amount of shared information for PnP-CoSMo to be effective, specifically =1/42=1/16 times the local information of the full-resolution case.
@@ 5.1.5 Sensitivity to misalignment of the reference
While PnP-CoSMo expects an aligned reference image at reconstruction time, the CR module should, in principle, correct for small misalignments. To test this, we simulated a 2∘ rotation in the reference image and performed reconstruction at =4 with and without the CR module, comparing also with PnP-CoSMo that used an aligned reference image. As shown in Fig. 6, although the reconstruction was sensitive to the misalignment in the absence of CR, it significantly improved with CR enabled.
@@ 5.1.6 Influence of the CR step size
The hyperparameter in Algorithm 1 controls the CR step size. Fig. 7 shows reconstruction quality as a function of acceleration and k-space noise level, where we observe that the optimal decreased with the amount and quality of k-space data. This makes sense as enforcing agreement with the k-space gradually becomes less beneficial, thereby contributing less to reconstruction quality.
5.2 Benchmark on the NYU DICOM data
@@
In the NYU benchmarking setup considered here, let the T1W-guided T2W reconstruction task be denoted as Task-1 and the T2W-guided T1W reconstruction task as Task-2. Fig. 8 shows evaluation metrics for PnP-CoSMo and the end-to-end baselines on these two tasks. We observe three main trends. First, the end-to-end methods generally dropped in performance in the out-of-distribution cases (0.05 throughout, with the exception of SSIM at =5 and HaarPsi at {2,4,5} for MoDL on Task-2). Second, PnP-CoSMo outperformed out-of-distribution models (0.05, except DISTS of MoDL and MCVarNet at =2 on Task-2). Since PnP-CoSMo does not depend on any task-specific k-space training data owing to both its plug-and-play design and because of the bi-directionality of the content/style model it uses, it is not susceptible to the distribution shift of this kind. Third, considering the in-distribution model performance, PnP-CoSMo largely outperformed MoDL (0.05 throughout, except SSIM at =2 and DISTS at {2,3,4} on Task-1 and DISTS at =2 on Task-2) and MTrans (0.05 throughout, except SSIM at {3,4} and DISTS at {2,3}), and was roughly comparable to MC-VarNet, despite having lenient training data requirements. As shown in Fig. 9a, PnP-CoSMo could resolve certain fine details even at the high acceleration of =5, which the baselines failed at. In the pathological case shown in Fig. 9b, PnP-CoSMo produced sharper reconstructions while preserving the structure of the edema better than (the in-distribution) MC-VarNet, despite lower metrics, when many features of edema were absent in the reference image.
5.3 Benchmark and ablation on the LUMC multi-coil data
@@
The LUMC datasets represent a more challenging problem with real-world data constraints. Additionally, in the light of the content capacity analysis (Section 5.1.2), we empirically found that the content capacity =1/4 was optimal for the LUMC datasets, compared to NYU DICOM data where =1 was the optimal content capacity. This reflects the lower effective content in LUMC image data and hence a greater difficulty of the guided reconstruction problem.
@@
Fig. 10 plots the benchmark metrics for the LUMC test sets, which can be summarized in the following three trends. First, pure image translation with MUNIT was worse compared to single-contrast reconstruction, especially at lower acceleration factors (comparing with PnP-CNN, 0.05 for all metrics and both datasets). Combining it with L2-regularized reconstruction in PROSIT improved SSIM and HaarPSI (0.05 for both datasets and accelerations), but not necessarily DISTS, suggesting that the the available complementary information was not fully utilized. PnP-CoSMo was consistently better than both MUNIT image translation and PROSIT (0.05 throughout for both cases). It also outperformed the conventional guided CS-STV (0.05 throughout, except SSIM at =4 in LUMC-COR). Hence, using both measured k-space and reference scan via PnP-CoSMo was more beneficial than (a) using either one of them and (b) combining both using hand-crafted priors, suggesting that our approach maximally exploits the complementary information. Second, both PnP-CNN and PnP-CoSMo performed similarly at lower acceleration, except in terms of SSIM and HaarPSI at =4 and =6 on LUMC-COR where PnP-CNN was slightly better. At higher acceleration, PnP-CoSMo outperformed PnP-CNN (0.05 for all metrics and both datasets except SSIM at =8 in LUMC-COR). Compared to PnP-CNN, PnP-CoSMo allowed up to 32.6% more acceleration for a given SSIM.222Based on linear interpolation of LUMC-TRA SSIM, PnP-CoSMo and PnP-CNN allowed =10 and =6.7, which is the maximum difference in , at median SSIM=0.906. Hence, a difference of 32.6%. Third, introducing PFT and CR into the ablated PnP-CoSMo improved the reconstructions (0.05 throughout in both cases except in DISTS for the latter at =10 in LUMC-TRA). The contribution of PFT was almost constant across , whereas that of CR decreased as increased, which was expected since CR depends on the measured k-space data to refine the content.
@@
5.4 Radiological evaluation
Figure 12 plots the radiological evaluation result. At =3, both algorithms produced diagnostic-quality reconstructions, a difference of 33.3% of k-space samples over the clinical reconstructions of =2. PnP-CNN was slightly better than PnP-CoSMo at in terms of sharpness, contrast, and the pathology criteria. At =5, PnP-CoSMo matched or exceeded PnP-CNN in terms of visual quality, although being worse in terms of pathology criteria. At =6, PnP-CNN sharply dropped to the ”non-diagnostic” level, whereas PnP-CoSMo was better, especially in terms of sharpness and contrast as well as in overall diagnostic quality. This trend is illustrated with a representative example in Figure 13. While the =3 reconstructions of PnP-CNN and PnP-CoSMo are of a similarly high quality, reconstruction artifacts (cyan arrows) become more pronounced in PnP-CNN reconstructions with increasing acceleration. At =6, compared to PnP-CoSMo, the PnP-CNN reconstruction has severe artifacts, blurrier and more distorted features (purple arrows), and a poorly visible hyperintense area of a brain metastasis (salmon-pink arrows).
@@ 6 Discussion
Optimal content capacity was defined as a quantity representing the amount of shared local information contained in a two-contrast dataset. In terms of this quantity, we analyzed the effect of reference image resolution on reconstruction quality, obtaining a lowerbound for PnP-CoSMo to be effective (Section 5.1.2). This quantity would depend on more fundamental factors such as MR sequence types, which we did not investigate here. While we limited our experiments to T1W and T2W sequences, PnP-CoSMo is, in principle, applicable to any pair of contrasts, with the image quality limited mainly by the amount of structural information shared between them. Future work could help empirically determine which sequence pairs are more amenable to guided reconstruction than others. A consequence of our MUNIT-based model design was the model’s lack of dependence on the uni-directionality of the guided reconstruction problem, which allowed greater generalizability by design (Section 5.2). It is, in principle, also possible to model the content and style for more than two contrasts. This would enable not only the generalizability of our approach to any combinatorial set of guided reconstruction tasks, but also better estimation of the target contrast’s content by utilizing multiple reference contrasts in the reconstruction. Given number of contrasts , content could be defined as the local structure underlying all these contrasts, while style would encode intra-contrast variations of each individual contrast. This model would then require an (unpaired) image-domain training dataset of the contrasts. An open question is how to effectively scale the model and training complexity to contrasts. The PnP-CoSMo algorithm can subsequently be extended to utilize multiple reference contrasts to guide a given target contrast. The specific question to investigate then would be – given reference contrasts (where ), how to aggregate their contents such that the prior structural information about the target image is maximized?
@@
Content discrepancy was characterized as a quantity representing the gap between the true content of the target image and our estimation of this content, which is another limiting factor in PnP-CoSMo reconstruction. In our MUNIT-based network design, the decoders model the (forward) generative process that maps the latent variables (i.e. content and style) to the observable variable (i.e. images). Ideally, the encoders must be a perfect inverse of this forward process. However, estimating the multi-channel contrast-independent content from a single image (using the content encoders) is a challenging (and perhaps ill-conditioned) inverse problem. Hence, some errors are to be expected in the content estimated from the reference image. Moreover, the reference image may contain unfavorable differences from the target image, such as misalignment or artifacts, which introduce more errors. Therefore, content discrepancy is expected to be a non-zero quantity. We proposed two complementary ways to minimize its effect on the reconstruction. While paired fine-tuning improves the model using a small amount of paired image-domain training data, content refinement (CR) aims to correct the remaining discrepancy during reconstruction using the measured k-space. A drawback of the CR module, however, is slower convergence (Section 5.1.3). For example, the run-time for PnP-CoSMo on the GPU was around 17 seconds for a slice of matrix size 349284, as compared to 6 seconds for CS-WT, 1.4 seconds for PnP-CNN, and less than 0.1 second for MC-VarNet. This latency issue could be mitigated by approximating the iterative process with an unrolled network design, although potentially at a cost of the generalizability offered by the plug-and-play design. The resulting trade-off between model latency and generalizability would be interesting to explore in the future.
@@
In terms of training data requirements, PnP-CoSMo can make use of the larger amounts of the image data available and can be applied to situations where end-to-end methods are infeasible since it does not rely on k-space training data. In situations where end-to-end methods are indeed feasible, PnP-CoSMo is comparable to state-of-the-art methods in terms of metrics and superior in terms of visual quality, especially at higher accelerations (Section 5.2 ). Moreover, unpaired image-domain pre-training of the content/style model boosts the practical applicability of PnP-CoSMo, e.g. in the case of LUMC-TRA where the T1W/T2W data imbalance was considerable. As indicated in our NYU DICOM benchmark, PnP-CoSMo performs similarly or better than state-of-the-art end-to-end methods. In our LUMC multi-coil benchmark, we limited our baselines to those algorithms that were feasible given the data constraint. As future work, a more comprehensive study should be conducted with a broader range of methods and evaluation criteria, e.g. comparing generalizability across accelerations and sampling patterns with multi-contrast unrolled networks, given a fixed budget of paired training data. Contrary to the training stage, we strictly assumed the spatial alignment of the reference and target images in the reconstruction stage. The CR module, which can implicitly correct minor registration errors, would likely break down at the typically observed levels of patient motion. A potential solution is to incorporate an online registration step to explicitly and efficiently correct for arbitrary inter-scan motion, thereby further improving practical applicability.
@@
In terms of reconstruction quality, we observed in Section 5.3 that the true added value of PnP-CoSMo was at the high acceleration factors (=8 and =10) where the k-space data is scarce and the reference information becomes more valuable. However, the risk of model hallucinations increases at these accelerations, raising concerns about the accuracy of the anatomy represented in the image. We hypothesize that as a consequence of content/style decomposition, it may be possible to leverage the contrast-independence property of the content to – (a) automatically detect local hallucinations in the reconstruction and (b) define a corrective process in combination with an adaptive sampling strategy to minimize it. This is yet another topic for future work. At lower accelerations, on the other hand, we observed that PnP-CoSMo often produced slightly lower metrics compared to PnP-CNN, e.g. in LUMC-COR. This was also observed in the radiological evaluation on LUMC-TRA samples. This is counter-intuitive at first glance since PnP-CoSMo has access to additional side information and should, in principle, perform at least as well as PnP-CNN. However, this can be explained by the fact that at lower accelerations, the content/style model becomes the main limiting factor for reconstruction quality. This is because unlike the CNN denoiser (Eq. (4)) of PnP-CNN, the content consistency operator (Eq. (15)) of PnP-CoSMo radically alters the image and hence, this operation is sensitive to the model’s overall performance. The model performance is in turn influenced by the training dataset and the network architecture. First, regarding data, the T1W scans in LUMC datasets had an in-plane resolution 1.3-2.5 times lower than that of the T2W scans, limiting the quality of the model (as indicated by its low optimal content capacity) and the overall value of guidance. In the long run, this may be solved by using 3D reference and target sequences of matched-resolution to optimize guidance. Second, regarding model design, correcting for content discrepancy using the CR module relies on computing gradients through the decoder (Eq. (17)). If is highly non-linear and the reference content is not sufficiently close to the optimum, CR may converge to a sub-optimal point leading to texture artifacts in the reconstruction, as seen in Fig. 11 at =4. In our proof-of-concept, we used the same general model architecture proposed in Huang et al. (2018). Better designs that are more conducive to CR may exist. Hence, future efforts must considerably focus on improving the content/style model using specialized architecture and regularization.

### v4

@@ 4.2 Implementation details
We used a similar general residual architecture for content/style encoders and decoders as in the original MUNIT paper (Xuan et al., 2022), except with an additional layer at the end of the content encoder to produce content maps of the specified number of channels. We used 2 or 4 content channels, depending on the content downsampling factor. The discriminators were implemented as multi-scale PatchGAN networks enabling them to locally assess the input images for realism at different scales. We used 1 scale in the BrainWeb simulations and 3 scales in the NYU and LUMC benchmark. We additionally conditioned the discriminators on foreground masks of the images to penalize background signal, and we used spectral normalization (Miyato et al., 2018) to stabilize the GAN training. For hyperparameters, we found that pre-training loss weight values and paired fine-tuning values worked generally well for the in vivo datasets. In the benchmarks, the step size of the CR module of PnP-CoSMo was tuned per acceleration for each dataset and the number of iterations was set to 200.
In the simulations, all content/style models were pre-trained for 200k and were not fine-tuned for the sake of simplicity, unless stated otherwise in Section 5.1. In both NYU and LUMC benchmarks, the content/style models were pre-trained for 400k iterations and fine-tuned for 50k iterations. We empirically found with a content capacity =1 of the content/style model as optimal for NYU dataset, whereas models with capacity =1/4 were optimal for the LUMC datasets. All software was implemented in PyTorch, and image registration was performed using Elastix (Klein et al., 2009). All training runs were performed on a compute node with an NVIDIA Quadro RTX 6000 GPU.
@@ 5 Results
5.1 Analyses on simulated MR data
@@ 5.1.2 Content capacity analysis
Table 4 compares PnP-CoSMo reconstruction quality across these configurations. PFT and CR were disabled here for simplicity. We observe two trends. First, the reconstruction quality generally decreases from left to right, eventually dropping below the conventional L1-wavelet CS reconstruction, suggesting a decrease in the amount of shared local information contained in the reference contrast. Second, the optimal content capacity of the model decreases with the reference resolution in accordance with the actual amount of this shared information. In other words, this optimal content capacity corresponds to the amount of contrast-independent structure the model discovers in the dataset, e.g. RefRes-1 dataset had 4 times as much shared structure (due to twice as large optimal content resolution) as RefRes-1/2 and RefRes-1/4 datasets and hence, the RefRes-1 dataset was significantly more effective in guided reconstruction. On the other hand, RefRes-1/2 and RefRes-1/4 datasets had similar levels of shared structure (i.e. =1/22=1/4 times the full-resolution case) despite the lower reference resolution in RefRes-1/4, suggesting a more complex relationship between the reference resolution and the shared content. In the RefRes-1/8 case, PnP-CoSMo dropped to a similar level of quality as conventional CS, suggesting a lowerbound on the amount of shared information for PnP-CoSMo to be effective, specifically =1/42=1/16 times the local information of the full-resolution case.
@@ 5.1.5 Sensitivity to misalignment of the reference
While PnP-CoSMo expects an aligned reference image at reconstruction time, the CR module should, in principle, correct for small misalignments. To test this, we simulated a 2∘ rotation in the reference image and performed reconstruction at =2 with and without the CR module, comparing also with PnP-CoSMo that used an aligned reference image. As shown in Fig. 6, although the reconstruction was sensitive to the misalignment in the absence of CR, it significantly improved with CR enabled.
@@ 5.1.6 Influence of the CR step size
The hyperparameter in Algorithm 1 controls the CR step size. Fig. 7 shows reconstruction quality as a function of acceleration and k-space noise level, where we observe that the optimal decreased with the amount and quality of k-space data. This was expected as enforcing agreement with the k-space gradually becomes less beneficial, thereby contributing less to reconstruction quality.
5.2 Benchmark on public NYU fastMRI DICOM data
@@
In the NYU benchmarking setup considered here, let the T1W-guided T2W reconstruction task be denoted as Task-1 and the T2W-guided T1W reconstruction task as Task-2. Fig. 8 shows evaluation metrics for PnP-CoSMo and the end-to-end baselines on these two guided reconstruction tasks. We observe three main trends. First, the end-to-end methods consistently dropped in performance on the out-of-distribution tasks (0.05 for all 3 models on both tasks at all 4 accelerations). Second, PnP-CoSMo outperformed all models on out-of-distribution tasks (0.05 throughout). Since PnP-CoSMo does not depend on any task-specific k-space training data owing to both its plug-and-play design and because of the bi-directionality of the content/style model it uses, it is not susceptible to the distribution shift of this kind. Third, considering the in-distribution model performance, PnP-CoSMo largely outperformed MoDL (0.05 throughout, except SSIM at =2 and DISTS at on Task-1, and DISTS at =2 on Task-2) and MTrans (0.05 throughout), and was roughly comparable to MC-VarNet (e.g. with 0.05 on Task-2 in terms of DISTS at =3 and SSIM at =4) while often outperforming it (0.05 throughout, except SSIM at and DISTS at on Task-1, and SSIM at and DISTS at =2 on Task-2). As shown in Figs. 9 for Task-2 and 10a for Task-1, PnP-CoSMo could resolve certain fine details even at the high acceleration of =5, which the end-to-end baselines failed at. In the pathological case shown in Fig. 10b, PnP-CoSMo produced sharper reconstructions while preserving the structure of the edema better than (the in-distribution) MC-VarNet, despite slightly lower metrics, when many features of the edema were absent in the reference image.
5.3 Benchmark and ablation on in-house LUMC multi-coil data
@@
The LUMC datasets represent a more challenging problem with real-world data constraints. Additionally, in the light of the content capacity analysis (Section 5.1.2), we must note that the content capacity =1/4 was optimal for the LUMC datasets, compared to NYU fastMRI DICOM data where =1 was the optimal content capacity, which reflects a lower effective content information in the LUMC images and hence a greater difficulty of the guided reconstruction problem.
@@
Fig. 11 plots the benchmark metrics for the LUMC test sets, which can be summarized in the following three trends. First, pure image translation with MUNIT was worse compared to single-contrast reconstruction, especially at lower acceleration factors (comparing with PnP-CNN, 0.05 for all metrics and both datasets). Combining it with L2-regularized reconstruction in PROSIT improved SSIM and HaarPSI (0.05 for both datasets and accelerations), but not necessarily DISTS, suggesting that the available complementary information was not fully utilized. PnP-CoSMo was consistently better than both MUNIT image translation and PROSIT (0.05 throughout for both cases). It also outperformed the conventional guided CS-STV (0.05 throughout, except SSIM at =4 in LUMC-COR). Hence, using both measured k-space and reference scan via PnP-CoSMo was more beneficial than (a) using either one of them and (b) combining both using hand-crafted priors, suggesting that our approach maximally exploits the complementary information. Second, both PnP-CNN and PnP-CoSMo performed similarly at lower acceleration, except in terms of SSIM and HaarPSI at =4 and =6 on LUMC-COR where PnP-CNN was slightly better. At higher acceleration, PnP-CoSMo outperformed PnP-CNN (0.05 for all metrics and both datasets except SSIM at =8 in LUMC-COR). Compared to PnP-CNN, PnP-CoSMo allowed up to 32.6% more acceleration for a given SSIM.222Based on linear interpolation of LUMC-TRA SSIM, PnP-CoSMo and PnP-CNN allowed =10 and =6.7, which is the maximum difference in , at median SSIM=0.906. Hence, a difference of 32.6%. Third, introducing PFT and CR into the ablated PnP-CoSMo improved the reconstructions (0.05 throughout in both cases except in DISTS for the latter at =10 in LUMC-TRA). The contribution of PFT was almost constant across , whereas that of CR decreased as increased, which was expected since CR depends on the measured k-space data to refine the content.
@@
5.4 Preliminary radiological evaluation
Fig. 13 plots the aggregated scores of our preliminary radiological evaluation and Fig. 14 shows a set of representative example reconstructions. At , PnP-CoSMo produced reconstructions of at least ”fairly diagnostic” level (Likert score 3) of image sharpness, artifacts, tissue contrast, and overall diagnostic quality, while enabling the detection of at least 80% of all metastases. In terms of artifacts, PnP-CoSMo was consistently superior, as seen also in Fig. 14 (blue arrows). In terms of tissue contrast, it matched or exceeded PnP-CNN at and produced sharper reconstructions at (purple arrows in Fig. 14). However, in terms of noise, it was equal or slightly worse. The may be explained by the difference in nature between the denoiser-type model of PnP-CNN and the more complex content/style model of PnP-CoSMo, potentially leading to noise-like grainy texture in the latter which also interferes with image sharpness and tissue contrast. In terms of the pathology criteria, while both algorithms scored similarly on the metastasis count, PnP-CoSMo had slightly better diagnostic quality at =6. The general difference in reconstruction quality at =6 can be seen in Fig. 14 where, compared to PnP-CoSMo, the PnP-CNN reconstruction shows severe reconstruction artifacts, blurrier and more distorted features (purple arrows), and a poorly visible hyperintense area of a brain metastasis (salmon-pink arrows).
In summary, both PnP-CoSMo and PnP-CNN allowed a maximum acceleration of =5 for at least fairly diagnostic overall quality, i.e. with 60% less k-space samples than the clinical reconstructions of =2. PnP-CoSMo was generally superior in terms of reconstruction artifacts and tissue contrast. Even at =6, the highest acceleration considered here, the strong multi-contrast prior in PnP-CoSMo enabled sharper reconstructions (0.67 higher mean Likert score) with greater tissue contrast (0.5 higher), less artifacts (0.67 higher), and higher overall diagnostic quality (0.17 higher) compared to the PnP-CNN baseline. The assessment by the two observers agreed with a Spearman’s rank correlation coefficient of 0.81 for the Likert-scale scores and 0.98 for the metastasis count.
@@ 6 Discussion
Optimal content capacity (Section 3.2.3) was defined as a quantity representing the amount of shared local information contained in a two-contrast dataset. In terms of this quantity, we analyzed the effect of reference image resolution on reconstruction quality, obtaining a lowerbound for PnP-CoSMo to be effective (Section 5.1.2), specifically a reference resolution 1/8 of the target image resolution, corresponding to =1/16 of the total spatial information in the target. This quantity would depend on more fundamental factors such as MR sequence types, which we did not investigate here. While we limited our experiments to T1W and T2W sequences, PnP-CoSMo is, in principle, applicable to any pair of contrasts, with the image quality limited mainly by the amount of structural information shared between them. Future work could help empirically determine which sequence pairs are more amenable to guided reconstruction than others. A consequence of our MUNIT-based model design was the model’s lack of dependence on the uni-directionality of the guided reconstruction problem, which allowed greater generalizability by design (Section 5.2). It is, in principle, also possible to model the content and style for more than two contrasts. This would enable not only the generalizability of our approach to any combinatorial set of guided reconstruction tasks, but also better estimation of the target contrast’s content by utilizing multiple reference contrasts in the reconstruction. Given number of contrasts , content could be defined as the local structure underlying all these contrasts, while style would encode intra-contrast variations of each individual contrast. This model would then require an (unpaired) image-domain training dataset of the contrasts. An open question is how to effectively scale the model and training complexity to contrasts. The PnP-CoSMo algorithm can subsequently be extended to utilize multiple reference contrasts to guide a given target contrast. The specific question to investigate then would be – given reference contrasts (where ), how to aggregate their contents such that the prior structural information about the target image is maximized?
@@
Content discrepancy (Section 3.3) was characterized as a quantity representing the gap between the true content of the target image and our estimation of this content, which is another limiting factor in PnP-CoSMo reconstruction. In our MUNIT-based network design, the decoders model the (forward) generative process that maps the latent variables (i.e. content and style) to the observable variable (i.e. images). Ideally, the encoders must be a perfect inverse of this forward process. However, estimating the multi-channel contrast-independent content from a single image (using the content encoders) is a challenging (and perhaps ill-conditioned) inverse problem. Hence, some errors are to be expected in the content estimated from the reference image. Moreover, the reference image may contain unfavorable differences from the target image, such as misalignment or artifacts, which introduce more errors. Therefore, content discrepancy is expected to be a non-zero quantity. We proposed two complementary ways to minimize its effect on the reconstruction. While paired fine-tuning improves the model using a small amount of paired image-domain training data, content refinement (CR) aims to correct the remaining discrepancy during reconstruction using the measured k-space. A drawback of the CR module, however, is slower convergence (Section 5.1.3). For example, the run-time for PnP-CoSMo on the GPU was around 17 seconds for a slice of matrix size 349284, as compared to 6 seconds for CS-WT, 1.4 seconds for PnP-CNN, and less than 0.1 second for MC-VarNet. This latency issue could be mitigated by approximating the iterative process with an unrolled network design, although potentially at a cost of the generalizability offered by the plug-and-play design. The resulting trade-off between model latency and generalizability would be interesting to explore in the future.
@@
In terms of training data requirements, PnP-CoSMo can make use of the larger amounts of the available image-domain data and can be applied to situations where end-to-end methods are infeasible. In situations where end-to-end methods are indeed feasible, such as in our NYU fastMRI benchmark (Section 5.2), PnP-CoSMo was often superior to state-of-the-art methods in terms of both metrics and visual quality, especially at higher accelerations, e.g. at =5 with a difference of 0.01 SSIM, 0.04 HaarPSI, and 0.02 DISTS on average over MC-VarNet in Task-1 and of 0.01 SSIM, 0.08 HaarPSI, and 0.04 DISTS in Task-2. Additionally, PnP-CoSMo is, by design, immune to cross-contrast distribution shifts as opposed to end-to-end methods, which was demonstrated, e.g. in comparison with MC-VarNet which suffered a difference of up to 0.10 SSIM, 0.17 HaarPSI, and 0.06 DISTS on average when applied on out-of-distribution tasks. Moreover, unpaired image-domain pre-training of the content/style model boosts the practical applicability of PnP-CoSMo, e.g. in the case of LUMC-TRA where the T1W/T2W data imbalance was considerable. In our LUMC multi-coil benchmark, we limited our baselines to those algorithms that were feasible given the data constraint. As future work, a more comprehensive study should be conducted with a broader range of methods and evaluation criteria, e.g. comparing generalizability across accelerations and sampling patterns with multi-contrast unrolled networks, given a fixed budget of paired training data. Contrary to the training stage, we strictly assumed the spatial alignment of the reference and target images in the reconstruction stage. The CR module, which can implicitly correct minor registration errors, would likely break down at the typically observed levels of patient motion. A potential solution is to incorporate an online registration step to explicitly and efficiently correct for arbitrary inter-scan motion, thereby further improving practical applicability.
@@
In terms of reconstruction quality on multi-coil data, we observed in Section 5.3 that the true added value of PnP-CoSMo was at the high acceleration factors () where the k-space data is scarce and the reference information becomes more valuable. For example, at =10 and compared to PnP-CNN, PnP-CoSMo showed an improvement of 0.04 SSIM, 0.08 HaarPSI, and 0.05 DISTS on LUMC-TRA and of 0.02 SSIM, 0.12 HaarPSI, and 0.09 DISTS on LUMC-COR. However, the risk of model hallucinations increases at these extreme accelerations, raising concerns about the accuracy of the anatomy represented in the image. We hypothesize that as a consequence of content/style decomposition, it may be possible to leverage the contrast-independence property of the content to – (a) automatically detect local hallucinations in the reconstruction and (b) define a corrective process in combination with an adaptive sampling strategy to minimize it. This is yet another topic for future work. At lower accelerations, on the other hand, we observed that PnP-CoSMo often produced slightly lower metrics compared to PnP-CNN, e.g. in LUMC-COR. This was also observed in our preliminary radiological evaluation on LUMC-TRA samples (Section 5.4), with PnP-CoSMo reconstructions at containing noise-like grainy texture which also interfered with criteria such as sharpness. At first glance, this is counter-intuitive since PnP-CoSMo has access to additional side information and should, in principle, perform at least as well as PnP-CNN. However, this can be explained by the fact that at lower accelerations, the quality of the content/style model becomes the main limiting factor for reconstruction quality. This is because unlike the CNN denoiser (Eq. (4)) of PnP-CNN, the content consistency operator (Eq. (15)) of PnP-CoSMo radically alters the image and hence, this operation is sensitive to the model’s overall performance. The model performance is in turn influenced by the training dataset and the network architecture. First, regarding data, the T1W scans in LUMC datasets had an in-plane resolution 1.3-2.5 times lower than that of the T2W scans, limiting the quality of the model (as indicated by its low optimal content capacity) and the overall value of guidance. In the long run, this may be solved by using 3D reference and target sequences of matched-resolution to optimize guidance. Second, regarding model design, correcting for content discrepancy using the CR module relies on computing gradients through the decoder (Eq. (17)). If is highly non-linear and the reference content is not sufficiently close to the optimum, CR may converge to a sub-optimal point leading to grainy texture in the reconstruction, e.g. as seen in Fig. 12 at =4. In our proof-of-concept, we used the same general model architecture proposed in Huang et al. (2018). Better designs that are more conducive to CR may be possible. Hence, future efforts will focus considerably on improving the content/style model using specialized architecture and regularization.
Finally, our preliminary radiological study demonstrated that (a) PnP-CoSMo was comparable with non-guided PnP-CNN in delivering at least fairly diagnostic overall quality (Likert score 3) reconstructions at =5, i.e. with 60% less k-space samples than the current clinical protocol of =2 and (b) PnP-CoSMo’s content/style modeling effectively leveraged reference contrast information enabling reconstructions with greater sharpness and tissue contrast, less artifacts, and a higher overall diagnostic quality at =6, with an improvement of 0.67, 0.5, 0.67, and 0.17 mean Likert scores, respectively. As a supplementary component to our LUMC multi-coil benchmark, this preliminary evaluation provides an early indication of potential benefit in the clinic, particularly time-saving during acquisition. Given the limited scale of our current evaluation, however, a large-scale clinical validation involving larger sample size and multiple radiologists must be conducted to robustly gauge the clinical utility of our approach. Such a clinical validation was beyond the scope of our current technical work and is, hence, a topic for a clinically-focused future work.